OpenStudy (anonymous):

did I do this right? (³√-125)³ simplifies to -125? or is that only if the second ^3 is inside ( )

OpenStudy (dumbcow):

no you are correct

OpenStudy (anonymous):

the cube root of -125 is -5 because -5 cubed is -125

OpenStudy (anonymous):

oh that makes sense.

OpenStudy (angela210793):

yea it simplifies only in -125 \[(\sqrt[n]{x})^n=x\]

OpenStudy (anonymous):

oh yes of course. i did not see the "cubed" part. you get what you started with good work!

myininaya (myininaya):

\[\sqrt[3]{(-5)^3}=-5=> (-5)^3=-125\]